In an internal combustion engine, air at atmospheric pressure and a temperature of about 11 ∘C is compressed in the cylinder by a piston to 1/8 of its original volume (compression ratio = 8.0).

Estimate the temperature of the compressed air, assuming the pressure reaches 44 atm .

In an internal combustion engine, air at atmospheric pressure and a temperature of about 17 ∘C is compressed in the cylinder by a piston to 1/8 of its original volume (compression ratio = 8.0).

Estimate the temperature of the compressed air, assuming the pressure reaches 46 atm .
Answer #1

Solution-

p1*V1/T1 = p2*V2/T2

So T2 = p2*V2*T1/(p1*V1) = 44*(1/8)*(273+11)/(1*1)

= 1562 K

so T2 =1562 - 273

= 1289 oC

b)

= 46*(1/8)*(273+17)/(1*1)

= 1667.5 K

so T2 =1667.5 - 273

=1394.5 oC